OpenStudy (mackenzie2013):

What is the value of the 11th term in the following geometric sequence? -5, -10, -20, -40 -5,120 -100 -180 -19,531,250

OpenStudy (***[isuru]***):

can u find the common ratio ?

OpenStudy (***[isuru]***):

if the common ratio is (r) and a_1 = -5 a_ 2 = -10 \[r = \frac{ a_{2} }{ a_{1}}\]

OpenStudy (mackenzie2013):

Isnt the common ratio 2?

OpenStudy (***[isuru]***):

after u find r .. u can use the general formula for a geometric progression \[T_{n} = a \ r^{n-1}\] where a = the first term = a_1= (-5) n = the nth term = in ur case 11 r = common ration = 2 ( yep u r correct! its 2) can u find it now ?

OpenStudy (***[isuru]***):

\[T_{11}= (-5) \times 2^{11 - 1}\]

OpenStudy (mackenzie2013):

So you take An= -5 x 2 (n-1)

OpenStudy (***[isuru]***):

yep!

OpenStudy (mackenzie2013):

So what N?

OpenStudy (***[isuru]***):

n = the n th term or the position of the term u want to find

OpenStudy (mackenzie2013):

So how do we find N?

OpenStudy (***[isuru]***):

u know the n... it;s the problem.. they ask the value of the term which is positioned at the 11th place in this sequence.. so... n = 11 got it :) ?

OpenStudy (mackenzie2013):

So AN = -5 x 2 (11-1=10) Then you take 10and add it too 2?

OpenStudy (***[isuru]***):

u have to find the 10th power of 2 \[2^{10}\] and multiply it by (-5)

OpenStudy (mackenzie2013):

so 20 x -5?

OpenStudy (***[isuru]***):

2x 10 = 2 multiply by 10 and its not what i mean 10 th power of 2 = 2 to the power 10 = 2x2x2x2x2x2x2x2x2x2 = ?

OpenStudy (mackenzie2013):

1,024

OpenStudy (***[isuru]***):

yep ... now multiply it by (-5)

OpenStudy (***[isuru]***):

got it :) ?

OpenStudy (mackenzie2013):

So 1,024 by -5?

OpenStudy (mackenzie2013):

So the answer is 5120?